WIENER RANDOM FIELDS WITH CONTROL MEASURES by Ljuban Dedić, Split, Croatia

Abstract. We deal with fields indexed by R^{n}. Using spectral measures, generated by filtration, we define stochastic integrals for them. We also represent some L_{2}- functionals by stochastic integrals and give some examples. Mathematics subject classification: 28C20 Keywords: Wiener fields, control measure, spectral measure, von Neumann algebrae.